# AICO: Feature Significance Tests for Supervised Learning

### v3

@@ 5.2 Classification
$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X))$ (16)
where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1) and the features $X=(X^{1},\ldots,X^{d})$ are as defined in Section 5.1. We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. The data are imbalanced, as is often the case in practice (e.g. credit scoring): about $10\%$ of the samples have $Y_{i}=1$. We train a single-layer neural network model $f$ of the conditional mean $\mathbb{E}(Y|X)$:
$f(x)=L\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right)$ (17)
where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters. The training procedure uses a standard cross-entropy criterion, $L_{2}$ regularization with weight $10^{-3}$ (approximately the optimal weight in the regression model training), Adam, a batch size of 32, and 500 epochs with early stopping. A validation set ($25\%$ of $I_{1}$) and a Bayesian approach are used to optimize the learning rate and the number of units, $K$. The optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The Area Under the ROC Curve on the training (and test) data exceeds $99\%$.
The testing procedure employs a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ consistent with the training criterion. We use the baseline and comparison feature specification (5.1). Table 4 reports the results for confidence levels $\alpha=0.01,0.05$. As in the regression case, the test has perfect power and avoids false discoveries. The test is generally robust to alternative specifications of the baseline (e.g. using the median instead of the mean to construct $v^{\ell}$, setting $v^{\ell}$ to the training mean also for the 1-hot feature $\ell=9$). [^13] The local baseline has perfect power but some false discoveries. The test is also robust to using an alternative score function $S(m,y)=-(m-y)^{2}$.
@@
Table 4 also reports the results for LOCO, CPT, HRT and Gaussian and deep knockoff procedures. AICO, CPT and LOCO have full power. LOCO, however, suffers from significant number of false discoveries (which cannot be addressed with a standard Bonferroni adjustment). HRT and knockoff procedures have less power and some false discoveries. In this comparison, AICO outperforms the alternatives. It performs as well as in the regression setting.
@@
Table 5 reports test results for varying test set sample sizes $N$. Power decreases faster with $N$ than in the regression case, especially for features $\ell=7,8$. False discoveries appear to be fairly insensitive to $N$.
@@
Table 6 provides a feature importance ranking according to the empirical median (6) for one of the 10 independent training/testing experiments that was randomly chosen. It also reports the endpoints of the randomized CI (7) for the population median $M$ at a $99\%$ level, and the standard (non-randomized) two-sided CI (9) along with coverage probabilities. Table 3 also provides the distribution of the randomized $p$-value.
Figure 6 shows the empirical median (6) for each of the ten training/testing experiments, for each of several methods including AICO. Across methods, the scores of the less important variables tend to be (somewhat) more volatile than in the regression setting. SHAP delivers inconsistent and highly unstable rankings, as in the regression case.
@@
Figure 7 reports average computation times across 10 experiments. As in the regression case, AICO achieves an unprecedented degree of computational efficiency. It is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is an order of magnitude more efficient than Gaussian knockoffs, the fastest screening procedure. It is more than three orders of magnitude faster than permutation importance, the fastest algorithmic feature importance measure.
@@ 6.1 Credit card default
This section complements our theoretical and numerical results with an empirical application to credit card payment data from 30,000 credit card holders of a Taiwanese card issuer, collected in October 2005.141414The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients The response variable is a binary indicator of payment default. There are 23 explanatory variables (see Table 7). The dataset was split into training (60%), validation (15%), and testing (25%) sets with class proportions maintained through stratified sampling. Categorical variables are one-hot encoded and continuous variables are standardized (mean 0 and unit variance) based on training set statistics.
@@
We implement an XGBoost (?) classifier using a cross-entropy loss function. We use the area under the ROC curve (AUC) on the validation set as a metric for hyper-parameter optimization. We employ the Optuna package (?), utilizing the Tree-Structured Parzen Estimator (?), with 250 trials. The search space includes the boosting algorithm, tree-specific settings such as tree depth and minimum child weight, and regularization terms. Class imbalance was mitigated using a scaling weight parameter, and training was capped at 1,000 boosting rounds with early stopping if the validation AUC did not improve for 50 consecutive rounds. The optimal hyper-parameters include a gradient tree boosting algorithm (?) with an approximation-based tree-building strategy; an L2 regularization parameter (lambda) of 0.2 and an L1 regularization parameter (alpha) of 0.001; a subsampling ratio of 100% (no subsampling); a column sampling ratio for each tree of 65%; a maximum tree depth of six levels; a minimum child weight of two; a learning rate (eta) of 0.05; a minimum loss reduction (gamma) of 0.00004; and a loss-guided tree growth policy.
@@
We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the same hyper-parameters obtained during the tuning phase. Each model was trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC. Ensemble predictions are obtained by averaging the outputs of all 100 models. The resulting ensemble model has AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set. (The results are robust with respect to the model. An ensemble of neural networks, for example, yields very similar results.)
@@
As in the classification example of Section 5.2, we use a score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to run the test. We continue to use the baseline and comparison feature specification (5.1), where $v^{\ell}_{i}$ is the mean of $\{X^{\ell}_{k}:k\in I_{1}\}$ in case feature $\ell$ is continuous. For a categorical variable we take the alternative mode $v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1})$ where $\mathcal{X}^{\ell}$ is the (marginal) support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the (empirical) probability mass function of $\{X^{\ell}_{k}:k\in I_{1}\}$. This specification yields values in the support of a categorical. Here, we treat a categorical variable as a single feature rather than treating the associated one-hot variables separately. Alternatively, if higher resolution is required, we can run the test for each individual level of a categorical.
@@
Table 8 summarizes the test results for $\alpha=1\%$ ($N=7,500$ throughout). The significant features, ranked according to the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE. Past delinquencies are significant predictors of future payment default. The more recent the payment record the stronger is the influence of timely payment. The actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant. Marital status, gender and age, but not education level, are significant predictors of default, even after controlling for past payment behavior. However, the influence of these variables is relatively small, especially for age (the associated medians are at least an order of magnitude smaller than the medians of the payment records). These results indicate that personal circumstances have a statistically significant influence on borrower behavior, beyond patterns in prior payment.
@@
We can further explore the role of a feature by running the test on subsets of the test set. For example, Table 9 reports the results for a test on the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$. While EDUCATION is not significant at the $1\%$ level, it is significant at the $5\%$ level (the distribution of the $p$-value is concentrated below $0.05$). Furthermore, EDUCATION is insignificant when running the test on the samples for which $\texttt{MARRIAGE}=\text{Married}$ (results available on request). This means educational level is a significant predictor of default for an unmarried borrower but not a married one.
@@
In the US, marital status, gender and age are protected characteristics under Fair Lending laws, so these variables are not used (they might not even be collected) to construct default probability estimates or credit scores for purposes of credit underwriting. Our results indicate, within the context of Taiwanese borrowers, that these attributes represent signals useful for risk prediction. All else equal, a score constructed without this set of variables might be less accurate.
@@ 6.2 Mortgage risk
We next consider AICO feature significance for a more complex classification model trained to predict the state of a mortgage loan in the next month given a history of a time-varying feature vector available at the beginning of the month. The 7 states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned. The analysis is based on an extensive panel dataset of mortgages from 4 ZIP codes in the greater Los Angeles area which are tracked monthly between 1/1994 and 12/2022 (this is a subset of the commercial CoreLogic Loan-Level Market Analytics dataset). The feature vector is of dimension $d=52$, and includes variables such as the current state, FICO credit score, loan balance, current interest rate, market interest rate, unemployment rate, etc., many of which are time-varying.
We use the data filtering procedure of ?, resulting in roughly 5 million loan-month transitions from 117,523 active loans. The train/validation/test splits are as follows: 1/1994 - 6/2009, 7/2009 - 12/2009, and 1/2010 - 12/2022. We adopt the set-sequence model of ? with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition data. We closely follow the training, validation and testing approach of ?.
@@ 7 Conclusion
This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning. Our randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies show strong performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, demonstrating its ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework lays the groundwork for a new standard of interpretability in machine learning—one that combines statistical grounding with scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
@@ Appendix B Panel data
Consider panel data $\{Z_{i,t}\}$ with $i=1,\ldots,m$ indexing units and $t=1,\ldots,T$ indexing time. We partition $\{(i,t):i\in[m],t\in[T]\}$ into two disjoint subsets, $I_{1}$ and $I_{2}$, such that each pair belongs to exactly one subset. This formulation allows for commonly used data-splitting schemes, such as temporal splits (distinguishing between in-sample and out-of-sample time periods), unit splits (assigning a unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both. In the case of an unbalanced panel, where units may be observed at different points in time, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
@@ B.1 Conditioning
Conditioning entails focusing on an appropriate subset $I\subseteq I_{2}$ when constructing the feature effect. For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to understand feature significance for a group of borrowers with certain features. If some samples in $I_{2}$ are dependent, as is often the case for panel data, we can choose $I$ as the subset of independent samples $(i,t)\in I_{2}$ (albeit at the cost of reducing the size of the test set). We can further partition $I$ by period $t$, to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by $t$. This approach leads to a sequence of tests, and a time series of $p$-values as well as importance scores and their CIs. The time-series behavior may provide useful insight.
For the purpose of constructing baseline and comparison features, the training set $I_{1}$ is also treated as a collection of $|I_{1}|$ samples–with unit and time information tied to each sample–fully harnessing panel data richness. For each $(i,t)\in I_{2}$, the baseline features can harness information from not only all training samples $I_{1}$ but also testing samples corresponding to the same unit $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$. This flexibility enables unit-specific and time-dependent baselines such as lagged feature value (e.g., last month’s feature value) and historical aggregation, enhancing the baseline’s meaningfulness.
@@ B.2 Grouping
An alternative grouping approach allows for joint testing across the time-series. The idea is to aggregate feature effects within a unit of dependence, so that dependence is contained within groups. We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predicted responses: $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method. Note that this differs slightly from the conditional approach. In the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only until the construction of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, unlike the conditional approach which includes the construction of the feature effect $\Delta_{i,t}$.
Then, we group the computed predictions along with the corresponding true responses by unit to compute the unit-level feature effect
$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i})$
where $S$ is a trajectory-level score function, $\underline{f}_{i}=\{f(\underline{X}_{i,t}):(i,t)\in I_{2}\}$, $\overline{f}_{i}=\{f(\overline{X}_{i,t}):(i,t)\in I_{2}\}$, and $Y_{i}=\{Y_{i,t}:(i,t)\in I_{2}\}$. Unlike previous single-period settings where the score function $S$ evaluates only one prediction against one true response, the trajectory-level score function evaluates the entire time series of predictions against the time series of true responses.
The trajectory-level score function can extend single-period score functions. An example is the negative mean squared error, where the average is across time steps. More complex score functions can harness rich temporal data, like correlation (Information Coefficient, IC), negative error in maximum drawdown (MDD): S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time series evaluations suitable for specific domain usage.
More broadly, the grouping approach extends beyond temporal dependence structures. For instance, for spatially dependent data, one could group testing samples by region or neighborhood, under the assumption that neighborhoods are (approximately) independent. Similarly, in settings with repeated measurements, grouping by the underlying entity or context may be effective.

### v4

@@ 5.2 Classification
$\mathbb{E}(Y|X)=\mathbb{P}(Y=1|X)=L(\mu(X)),$ (16)
where $L(u)=1/(1+e^{u})$, the function $\mu$ is given by (5.1), and the features $X=(X^{1},\ldots,X^{d})$ are defined as in Section 5.1. We generate a training set $I_{1}$ of $10^{6}$ independent samples $(X_{i},Y_{i})$ and a test set $I_{2}$ of $5\cdot 10^{5}$ independent samples. The data are imbalanced, as is common in practice (e.g., credit scoring): approximately 10% of the samples satisfy $Y_{i}=1$.
We train a single-layer neural network $f$ to estimate the conditional mean $\mathbb{E}(Y|X)$, specified as
$f(x)=L\!\left(b_{0}+\sum_{k=1}^{K}b_{k}\psi(a_{0,k}+a_{k}^{\top}x)\right),$ (17)
where $\psi(y)=1/(1+e^{-y})$, $K$ is the number of hidden units, and $b_{0},b_{k},a_{0,k}\in\mathbb{R}$ and $a_{k}\in\mathbb{R}^{d}$ are parameters. Training uses a standard cross-entropy loss with $L_{2}$ regularization (weight $10^{-3}$, approximately optimal in the regression setting), the Adam optimizer, a batch size of 32, and 500 epochs with early stopping. A validation set (25% of $I_{1}$) and a Bayesian search procedure are used to tune the learning rate and number of units $K$. The optimal learning rate is $5\cdot 10^{-4}$ and the optimal $K=300$. The resulting model achieves an Area Under the ROC Curve exceeding 99% on both the training and test data.
The testing procedure employs the score function $S(m,y)=y\log m+(1-y)\log(1-m)$, consistent with the training criterion. We use the baseline and comparison feature specification (5.1). Table 4 reports the results for confidence levels $\alpha=0.01$ and $0.05$. As in the regression case, the test achieves perfect power and produces no false discoveries. It remains robust to alternative baseline specifications—for example, using the median instead of the mean to construct $v^{\ell}$, or setting $v^{\ell}$ to the training mean even for the one-hot feature $\ell=9$. [^11] The local baseline attains perfect power but yields some false discoveries. The test is also robust to alternative score functions, such as $S(m,y)=-(m-y)^{2}$.
@@
Table 4 also reports results for LOCO, CPT, HRT, and the Gaussian and deep knockoff procedures. AICO, CPT, and LOCO all achieve full power. LOCO, however, produces a substantial number of false discoveries that cannot be corrected by a standard Bonferroni adjustment. HRT and the knockoff approaches exhibit lower power and some false discoveries. Overall, AICO outperforms all alternatives, matching its performance in the regression setting.
@@
Table 5 reports test results for varying test-set sample sizes $N$. Power declines more rapidly with smaller $N$ than in the regression case, particularly for features $\ell=7$ and $\ell=8$. The incidence of false discoveries appears largely insensitive to $N$.
@@
Table 6 presents a feature-importance ranking based on the empirical median (6) for one randomly chosen training/testing experiment out of ten. It also reports the endpoints of the randomized confidence interval (7) for the population median $M$ at the $99\%$ level, together with the standard (non-randomized) two-sided interval (9) and their coverage probabilities. Table 3 further shows the distribution of the randomized $p$-values.
Figure 6 displays the empirical median (6) across the ten independent training/testing experiments for several methods, including AICO. Across methods, the scores of the less important variables tend to be somewhat more volatile than in the regression setting. As before, SHAP yields inconsistent and highly unstable rankings across experiments.
@@
Figure 7 reports average computation times across ten experiments. As in the regression case, AICO achieves exceptional computational efficiency. It is more than three orders of magnitude faster than LOCO, the next most efficient testing procedure. AICO is roughly an order of magnitude faster than Gaussian knockoffs, the fastest screening approach, and more than three orders of magnitude faster than permutation importance, the most efficient algorithmic feature-importance method.
@@ 6.1 Credit card default
This section complements the theoretical and numerical results with an empirical application to credit-card payment data from 30,000 card holders of a Taiwanese issuer, collected in October 2005.121212The dataset is available at the UCI Machine Learning Repository: https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients. The response variable is a binary indicator of payment default, and there are 23 explanatory variables (see Table 7). The dataset is split into training (60%), validation (15%), and testing (25%) subsets, maintaining class proportions through stratified sampling. Categorical variables are one-hot encoded, and continuous variables are standardized to mean 0 and unit variance using training-set statistics.
@@
We implement an XGBoost classifier (?) using a cross-entropy loss function. Model performance is evaluated by the area under the ROC curve (AUC) on the validation set, which serves as the criterion for hyper-parameter optimization. We use the Optuna package (?) with the Tree-Structured Parzen Estimator (?), conducting 250 trials. The search space includes the boosting algorithm, tree-specific parameters such as maximum depth and minimum child weight, and regularization terms. Class imbalance is mitigated using a scaling-weight parameter, and training is capped at 1,000 boosting rounds with early stopping after 50 rounds without improvement in validation AUC. The optimal configuration employs gradient tree boosting (?) with an approximation-based tree-building strategy, an $L_{2}$ regularization parameter (lambda) of 0.2, an $L_{1}$ regularization parameter (alpha) of 0.001, a subsampling ratio of 100%, a column-sampling ratio per tree of 65%, a maximum tree depth of six, a minimum child weight of two, a learning rate (eta) of 0.05, a minimum loss reduction (gamma) of 0.00004, and a loss-guided tree-growth policy.
@@
We train an ensemble of 100 XGBoost models on 100 bootstrap samples of the training data, all using the tuned hyper-parameters. Each model is trained for up to 1,000 boosting rounds, with early stopping triggered after 50 rounds of no improvement in the validation AUC. Ensemble predictions are obtained by averaging the outputs of all 100 models. The resulting ensemble achieves AUCs of $0.837$ on the training set, $0.782$ on the validation set, and $0.786$ on the test set. The results are robust to the model choice; for example, an ensemble of neural networks yields nearly identical outcomes.
As in the classification example of Section 5.2, we use the score function $S(m,y)=y\log m+(1-y)\log(1-m)$ to conduct the test. We continue to apply the baseline and comparison feature specification (5.1), where $v_{i}^{\ell}$ is the mean of $\{X_{k}^{\ell}:k\in I_{1}\}$ when feature $\ell$ is continuous. For categorical variables, we use the alternative mode
$v_{i}^{\ell}=\arg\max_{x\in\mathcal{X}^{\ell}\setminus\{X_{i}^{\ell}\}}p^{\ell}(x|I_{1}),$
@@
where $\mathcal{X}^{\ell}$ denotes the marginal support of $X^{\ell}$ and $p^{\ell}(\cdot|I_{1})$ is the empirical probability mass function of $\{X_{k}^{\ell}:k\in I_{1}\}$. This specification ensures that the baseline value lies within the categorical support. We treat each categorical variable as a single feature rather than analyzing its associated one-hot encodings separately. If higher resolution is desired, the test can alternatively be run for each individual level of a categorical variable.
@@
Table 8 summarizes the test results for $\alpha=1\%$ (N=7,500N=7{,}500 throughout). The significant features, ranked by the empirical median (6), are PAY_0 through PAY_6, MARRIAGE, SEX, and AGE. Past delinquencies are strong predictors of future payment default, with more recent payment records exerting greater influence. Actual monthly bill and payment amounts (e.g., PAY_AMT and BILL_AMT) are not significant. Marital status, gender, and age—but not education level—are significant predictors of default even after controlling for past payment behavior. However, their influence is relatively small, especially for age: the corresponding medians are at least an order of magnitude smaller than those for the payment-history features. These findings suggest that personal circumstances exert a statistically significant, but comparatively modest, influence on borrower behavior beyond patterns in prior payment.
@@
We can further explore the role of individual features by running the test on subsets of the test set. For instance, Table 9 reports results for the 4,029 (out of 7,500) samples for which $\texttt{MARRIAGE}=\text{Single}$. While EDUCATION is not significant at the $1\%$ level, it becomes significant at the $5\%$ level, as the distribution of its $p$-value is concentrated below $0.05$. When the test is restricted to samples with $\texttt{MARRIAGE}=\text{Married}$ (results available upon request), EDUCATION is no longer significant. These findings indicate that educational level is a significant predictor of default for unmarried borrowers but not for married ones.
@@
In the United States, marital status, gender, and age are protected characteristics under Fair Lending laws and therefore are not used—indeed, often not even collected—for estimating default probabilities or credit scores in credit underwriting. Our findings suggest that, within the context of Taiwanese borrowers, these attributes carry predictive information relevant to risk assessment. All else equal, a score constructed without these variables may be less accurate.
@@ 6.2 Mortgage risk
We next examine AICO feature significance in a more complex classification setting, predicting the state of a mortgage loan in the next month given a history of time-varying features observed at the beginning of the month. The seven states are Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent, Foreclosure, Paid Off, and Real Estate Owned. The analysis is based on an extensive panel dataset of mortgages from four ZIP codes in the greater Los Angeles area, tracked monthly from January 1994 through December 2022. This dataset constitutes a subset of the CoreLogic Loan-Level Market Analytics data. The feature vector has dimension $d=52$ and includes variables such as current loan state, FICO credit score, loan balance, current interest rate, market interest rate, and local unemployment rate, many of which are time-varying.
We apply the data-filtering procedure of ?, resulting in roughly five million loan-month transitions from 117,523 active loans. The training, validation, and test splits are 1/1994–6/2009, 7/2009–12/2009, and 1/2010–12/2022, respectively. We adopt the set–sequence model of ? with LongConv (?) for the sequence layer, which represents the current state of the art for mortgage transition modeling. Training, validation, and testing procedures follow closely the approach of ?.
@@ 7 Conclusion
This paper develops a simple and computationally efficient framework for assessing feature significance in supervised learning. The proposed randomized sign test delivers exact $p$-values and confidence intervals for effect size under minimal assumptions, avoiding the limitations of retraining- and surrogate-model-based approaches while remaining robust to noise and feature correlation. Empirical studies demonstrate strong statistical performance on synthetic data and practical value in real-world applications such as credit scoring and mortgage risk modeling, illustrating the method’s ability to inform high-stakes decisions with legal, economic, and regulatory implications. More broadly, the framework establishes a foundation for a new standard of interpretability in machine learning—one that combines statistical rigor with computational scalability and directly supports trustworthy, evidence-based decision-making in science, industry, and policy.
@@ Appendix B Panel data
Consider panel data $\{Z_{i,t}\}$, where $i=1,\ldots,m$ indexes units and $t=1,\ldots,T$ indexes time. We partition $\{(i,t):i\in[m],t\in[T]\}$ into two disjoint subsets, $I_{1}$ and $I_{2}$, such that each pair belongs to exactly one subset. This setup accommodates common data-splitting schemes, including temporal splits (distinguishing between in-sample and out-of-sample time periods), unit-based splits (assigning each unit’s entire time series to either $I_{1}$ or $I_{2}$), or combinations of both. For an unbalanced panel, in which units may be observed at different times, only observed pairs $(i,t)$ are assigned to $I_{1}$ or $I_{2}$.
@@ B.1 Conditioning
Conditioning entails focusing on a suitable subset $I\subseteq I_{2}$ when constructing the feature effect. For example, in the credit card analysis of Section 6.1, we condition on card applicants who are married to assess feature significance within a specific borrower subgroup. If some samples in $I_{2}$ are dependent—as is common in panel data—we can restrict attention to the subset $I$ of independent samples $(i,t)\in I_{2}$, albeit at the cost of a smaller test set. We can further partition $I$ by period $t$ to construct feature effects $\{\Delta_{\cdot,t}:(\cdot,t)\in I\}$ indexed by time. This yields a sequence of tests and a time series of $p$-values, importance scores, and confidence intervals whose temporal dynamics may provide additional insight.
For constructing baseline and comparison features, the training set $I_{1}$ is treated as a collection of $|I_{1}|$ samples—with unit and time information tied to each—fully exploiting the richness of the panel structure. For each $(i,t)\in I_{2}$, the baseline features can draw information not only from all training samples $I_{1}$ but also from testing samples corresponding to the same unit, $\{X_{i,\cdot}:(i,\cdot)\in I_{2}\}$. This flexibility enables unit-specific and time-dependent baselines, such as lagged feature values (e.g., the previous month’s feature value) or historical aggregations, thereby enhancing the contextual relevance of the baseline.
@@ B.2 Grouping
An alternative grouping approach enables joint testing across time series. The idea is to aggregate feature effects within a unit of dependence so that dependence is contained within groups. We treat the testing set $I_{2}$ as a collection of $|I_{2}|$ samples and compute the baseline and treatment predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$ for $(i,t)\in I_{2}$, where $\underline{X}_{i,t}$ and $\overline{X}_{i,t}$ are baseline and treatment feature values constructed as in the conditional method. This differs slightly from the conditioning approach: in the grouping approach, $I_{2}$ is treated as a collection of $|I_{2}|$ samples only up to the computation of the baseline and comparison predictions $f(\underline{X}_{i,t})$ and $f(\overline{X}_{i,t})$, whereas the conditional approach proceeds to construct the feature effect $\Delta_{i,t}$ at the individual-sample level.
Next, we group the computed predictions and corresponding true responses by unit to obtain the unit-level feature effect
$\Delta_{i}=S(\overline{f}_{i},Y_{i})-S(\underline{f}_{i},Y_{i}),$
where $S$ is a trajectory-level score function, $\underline{f}_{i}=\{f(\underline{X}_{i,t}):(i,t)\in I_{2}\}$, $\overline{f}_{i}=\{f(\overline{X}_{i,t}):(i,t)\in I_{2}\}$, and $Y_{i}=\{Y_{i,t}:(i,t)\in I_{2}\}$. Unlike the single-period setting, where $S$ evaluates one prediction against one true response, the trajectory-level score function evaluates the entire sequence of predictions against the corresponding sequence of true responses.
Trajectory-level score functions can extend familiar single-period metrics. For instance, one may use the negative mean squared error, averaging across time steps. More complex specifications can exploit richer temporal structure, such as correlation (Information Coefficient, IC), the negative error in maximum drawdown (MDD), S(m,y)=|MDD(y)MDD(m)|S(m,y)=-|MDD(y)-MDD(m)|, or other time-series evaluations tailored to specific domains.
More broadly, the grouping approach extends beyond temporal dependence structures. For spatially dependent data, one may group testing samples by region or neighborhood, under the assumption that neighborhoods are approximately independent. Likewise, for repeated-measures data, grouping by the underlying entity or experimental context can provide an effective way to respect within-group dependence while maintaining valid inference across groups.